A blonde hair was used to form one end of an air wedge with two glass plates. The glass plates were in contact at the other end. When light of wavelength 460 nm was reflected from the air wedge, the 320th dark fringe coincided with the position of the hair. What was the thickness of the hair?
1
Expert's answer
2020-06-17T09:39:27-0400

As per the question,

The wavelength of the light "(\lambda)=460nm"

n =320 dark fringe

Let the thickness be e,

"e=\frac{m\lambda }{2}"

"\Rightarrow e=\frac{320\times 460\times 10^{-9}m}{2}"


"\Rightarrow e=73.6\times 10^{-6}m"

"\Rightarrow e =73.6\mu m"